# A Survey on Dynamic Neural Networks: from Computer Vision to Multi-modal Sensor Fusion

### v1

@@ 4 Discussion
Stability and convergence. Incorporating dynamic behavior within a network often translates in the need of performing discrete decisions. The training of discrete decision modules tends to lead to instability during the training process, hindering the convergence toward an optimal result or in some cases destabilizing the whole process. Furthermore, the discretization sometimes induces different parts of the network to learn from different number of subsamples related to the specific task, resulting in a poor learning on certain subtasks.
@@
Benchmarks and metrics. The lack of standardized metrics and benchmarks tailored to Dynamic Neural Networks complicates the comparison among different methods. Conventional metrics (e.g., accuracy, FLOPs) often fail to adequately compare the adaptive nature of these model.
Robustness. Robustness to adversarial attacks is an underexplored area. Preliminary studies (Zhang et al., 2023, Pan et al., 2023) provide some insights but are limited in scope. The full extent of possible attacks on any Dynamic Neural Network has yet to be explored. Furthermore, there are no studies comparing the robustness to attacks of these networks to static networks.
@@
Architecture agnosticism. Future research should prioritize broader diffusion of these techniques across diverse architectures. A crucial step is enabling any static pretrained network architecture to be converted to a dynamic model with minimal effort. For example, Bolya et al. (2023) and Seol et al. (2023) demonstrate approaches that introduce dynamicity with no extra parameters added to the original pretrained ViT, no need for model fine-tuning and a contained drop in accuracy with respect to the original model. This enables the solution to be plug-and-play to all ViTs classification models after the training phase with little additional effort.
@@
Sensor Fusion. When it comes to Dynamic Neural Networks in the context of Sensor Fusion, there are many unexplored areas. The temporal aspect has been largely explored (Sabet et al., 2022, Yang et al., 2024, Chen et al., 2023), and Dynamic Sensor Fusion could benefit from the ability to adaptively handle asynchronous or variable-rate data streams. Furthermore, different sensors produce data at different levels of detail, also depending on the context. Introducing dynamically the option of adapting the granularity of fused representations could help both the efficiency and the performance. A fully modular solution should also be explored, where components of a network can be added and removed according to sensor availability, and the network can orchestrate among both sensors available and sensors needed for the task. Finally, ensuring safety and trust in multi-sensor systems is a key element in certain fields, like autonomous driving, and it requires both some understanding and robustness towards unknown events.
@@ 5 Conclusions
This survey focuses on Dynamic Neural Networks techniques for Computer Vision and their applications in Sensor Fusion. We presented a taxonomy based on a logical division of the techniques according to the parts of the networks that are dynamically influenced by the input.
All models that constrained the network output position with respect to the input complexity fall under the Early Exits Section, where we presented the most important architectural solutions, relevant methods for improving the training, and the application of these methods to specific tasks and use cases. Then we proceeded with the review of Dynamic Routing techniques, where the computational path is created dynamically, either by constructing it on the fly or routing the computation to different model blocks. Also, in this case, a we present the application of these methods to various tasks. Lastly, in the Token Skimming section, we focused on Transformer-based models that either drop certain tokens along the block of the network (temporarily or permanently) or merge them together, based on an estimated token contribution to the desired task. The last part is dedicated to other methods and the applications of skimming to tasks that differ from image classification.
@@ Acknowledgements
[…]
Data availability The authors declare that all data supporting the findings of this survey are publicly available in the cited literature, as referenced throughout the manuscript.

### v3

@@ 4 Discussion
Stability and convergence. Incorporating dynamic behavior within a network often translates in the need of performing discrete decisions. The training of discrete decision modules tends to lead to instability during the training process, hindering the convergence toward an optimal result or in some cases destabilizing the whole process. Furthermore, the discretization sometimes induces different parts of the network to learn from different number of subsamples related to the specific task, resulting in a poor learning on certain subtasks. Exploring techniques that ensure that all network paths are adequately trained is a task that require further investigation.
@@
Benchmarks and metrics. The lack of standardized metrics and benchmarks tailored to Dynamic Neural Networks complicates the comparison among different methods. Conventional metrics (e.g., accuracy, FLOPs) are often not sufficient to capture the adaptive nature of these models. Currently, the literature relies on comparisons within the most similar architecture, making a broader comparison among different approaches hard to make.
Robustness. Robustness to adversarial attacks is an underexplored area. Preliminary studies (Zhang et al., 2023a; Pan et al., 2023) provide some insights but are limited in scope. The full extent of possible attacks on any Dynamic Neural Network has yet to be explored. Furthermore, there are no studies comparing the robustness to attacks of these networks to static networks.
@@
The field of Dynamic Neural Networks has come a long way and many possible direction have arisen for future directions.
Benchmark To enable fair cross-study analysis, the field would highly benefit from a standardized benchmark. As previously discussed, cross technique comparison is currently a challenge. We argue that such a benchmark must consistently evaluate not only accuracy but also inference speed, the variability of computations in relation to input complexity, and robustness in edge cases (e.g., small objects or targets located in the peripheral regions of the image).
Architecture agnosticism. Future research should prioritize broader diffusion of these techniques across diverse architectures. A crucial step is enabling any static pretrained network architecture to be converted to a dynamic model with minimal effort. For example, Bolya et al. (2023) and Seol et al. (2023) demonstrate approaches that introduce dynamicity with no extra parameters added to the original pretrained ViT, no need for model fine-tuning and a contained drop in accuracy with respect to the original model. This enables the solution to be plug-and-play to all ViTs classification models after the training phase with little additional effort. We argue that the development of post-training solutions that are architecture agnostic is to be preferred, mirroring the success of static efficiency methods like quantization and pruning.
@@
Sensor Fusion. When it comes to Dynamic Neural Networks in the context of Sensor Fusion, there are many unexplored areas. The temporal aspect has been largely explored (Sabet et al., 2022; Yang et al., 2024a; Chen et al., 2023a), and Dynamic Sensor Fusion could benefit from the ability to adaptively handle asynchronous or variable-rate data streams. Furthermore, different sensors produce data at different levels of detail, also depending on the context. Introducing dynamically the option of adapting the granularity of fused representations could help both the efficiency and the performance. A fully modular solution should also be explored, where components of a network can be added and removed according to sensor availability, and the network can orchestrate among both sensors available and sensors needed for the task. Finally, ensuring safety and trust in multi-sensor systems is a key element in certain fields, like autonomous driving, and it requires both some understanding and robustness towards unknown events.
@@ 5 Conclusions
This survey provides a comprehensive analysis of 161 Dynamic Neural Network methodologies for Computer Vision and their applications in Sensor Fusion. We propose a taxonomy categorizing the techniques according to their adaptive network component: Early Exits (output-level), Dynamic Routing (path-level), and Token Skimming (input-level). Furthermore, we examine the specific advantages of these architectures for multi-modal tasks and supplement this review with a comprehensive, up-to-date repository of implementations to facilitate future research. All models that constrained the network output position with respect to the input complexity fall under the Early Exits Section, where we presented the most important architectural solutions, relevant methods for improving the training, and the application of these methods to specific tasks and use cases. Then we proceeded with the review of Dynamic Routing techniques, where the computational path is created dynamically, either by constructing it on the fly or routing the computation to different model blocks. Also, in this case, a we present the application of these methods to various tasks. Lastly, in the Token Skimming section, we focused on Transformer-based models that either drop certain tokens along the block of the network (temporarily or permanently) or merge them together, based on an estimated token contribution to the desired task. The last part is dedicated to other methods and the applications of skimming to tasks that differ from image classification.
@@
Furthermore, we provided a discussion on the current limitations and future directions. A considerable highlight is that Token Skimming allows for direct comparison on the ImageNet benchmark, where SaiT currently holds the best overall accuracy, whereas Early Exits and Dynamic Routing still lack a shared standard for consistent evaluation.
@@ Acknowledgements
Data availability
The authors declare that all data supporting the findings of this survey are publicly available in the cited literature, as referenced throughout the manuscript.